OpenStudy (anonymous):

80=H exp[-10.29(1/50-1/35)] solve for H

OpenStudy (anonymous):

please show steps

OpenStudy (anonymous):

First solve for (1/50-1/35) and it is equal to -3/350 or -0.00857

OpenStudy (anonymous):

ok. its the exp part i cant figure out

OpenStudy (anonymous):

Now multiply this with the term -10.29 which is equal to .0882

OpenStudy (anonymous):

Okay! now we have to take log on both of the sides so after taking log we will get log(80)=logH^(.0882) or we can write it as log(80)=(.0882)*log(H)

OpenStudy (anonymous):

Now we will plug the value of log(80) which is equal to 1.90

OpenStudy (anonymous):

so we will get 1.90/.0082= loh(H)

OpenStudy (anonymous):

Sorry it is 1.90/.0882

OpenStudy (anonymous):

So we will get 21.54=log(h)

OpenStudy (anonymous):

So H=10^21.54

OpenStudy (anonymous):

Is this clear so far?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

You are most welcome!:)